Wabash Products uses a job costing system and applies overhead based on direct labor cost. Last year, manufacturing overhead was expected to be $2,829,000 and direct labor cost was estimated to be $1,150,000 (40,600 direct-labor hours). Actual manufacturing overhead amounted to $2,958,000 and direct labor cost was $1,250,000 (42,800 direct-labor hours).

Wabash allocates an over- or underapplied overhead to the three accounts—Work-in-Process Inventory, Finished Goods Inventory, and Cost of Goods Sold—based on those account balances. At the end of the year, the total amount in the three accounts (Work-in-Process Inventory, Finished Goods Inventory, and Cost of Goods Sold) was $5.90 million before any allocation. As a part of the process, you learn that Wabash allocated $102,960 to Cost of Goods Sold.

Required:

What will Wabash report as Cost of Goods Sold for the year?
Answer #1

Overhead rate = 2829000*100/1150000 = 246% of direct labor cost

Overhead applied = 1250000*2.46 = 3075000

Actual overhead = 2958000

Over applied overhead = 3075000-2958000 = 117000

Over applied allocated to Cost of goods sold = 102960/5900000*117000 = 2042

Cost of goods sold for the year = 102960-2042 = 100918